# Dance Any Beat: Blending Beats with Visuals in Dance Video Generation

### v1

@@ Abstract
The task of generating dance from music is crucial, yet current methods, which mainly produce joint sequences, lead to outputs that lack intuitiveness and complicate data collection due to the necessity for precise joint annotations. We introduce a Dance Any Beat Diffusion model, namely DabFusion, that employs music as a conditional input to directly create dance videos from still images, utilizing conditional image-to-video generation principles. This approach pioneers the use of music as a conditioning factor in image-to-video synthesis. Our method unfolds in two stages: training an auto-encoder to predict latent optical flow between reference and driving frames, eliminating the need for joint annotation, and training a U-Net-based diffusion model to produce these latent optical flows guided by music rhythm encoded by CLAP. Although capable of producing high-quality dance videos, the baseline model struggles with rhythm alignment. We enhance the model by adding beat information, improving synchronization. We introduce a 2D motion-music alignment score (2D-MM Align) for quantitative assessment. Evaluated on the AIST++ dataset, our enhanced model shows marked improvements in 2D-MM Align score and established metrics. Video results can be found on our project page: https://DabFusion.github.io.
@@ 1 Introduction
Music-to-dance generation offers significant potential for automating choreography, benefiting various real-world applications. Existing methods [26, 27, 47, 35] focus on generating joint sequences, which, despite being effective, often produce outputs that are less intuitive for users and require precise joint annotations during data collection. This complicates the process. The field of conditional image-to-video (I2V) generation—where videos are produced from a single still image guided by textual descriptions [54, 39, 11], driving videos [6, 65, 55], or pose sequences [38, 45, 44, 68]—offers a way to address these challenges. However, using music as a driving condition remains underexplored, presenting an opportunity for innovation. In this paper, we propose the Dance Any Beat Diffusion model (DabFusion), which generates dance videos from still images and music, enhancing intuitiveness and eliminating the need for joint annotations. DabFusion also pioneers the use of music as a condition in image-to-video synthesis.
DabFusion is highly versatile, capable of generating dance videos using the AIST++ dataset and animating individuals in new scenarios. As shown in Figure 1, we animate previously unseen individuals by merging their images with a background from the AIST++ dataset, bringing them to life through dance. This capability makes DabFusion practical for real-world applications, such as aiding dancers and choreographers in learning and practice, fostering engagement on social media, and enabling users without formal dance training to create personalized dance content. When generating dance videos using the AIST++ dataset, as illustrated in Figure 2, DabFusion excels in producing varied dance styles with different dancers from multiple perspectives and diverse initial poses. This flexibility enhances the intuitiveness and applicability of the generated content.
Inspired by recent I2V works [32, 44, 57], our model’s training process is structured into two phases. First, a latent flow auto-encoder is trained unsupervisedly to estimate the latent optical flow between reference and driving frames in a video, aiding in warping the reference frame to generate movement. Then, a U-Net based diffusion model generates latent flows guided by music and a starting image. For music information extraction, we use CLAP [59], the largest audio representation foundation model, to facilitate high-quality, music-aligned dance video generation.
A critical aspect of dance video evaluation is synchronization between the dancer’s movements and the music’s beat, an area needing improvement in our initial model. To address this, we integrate Librosa [30], a tool for audio signal analysis and beat extraction, with CLAP’s output. This enhancement improves alignment between dance poses and music beats, offering a more coherent and engaging viewing experience. To quantitatively assess this alignment, we introduce the 2D motion-music alignment (2D-MM Align) score, inspired by the Beat Alignment Score [27] for 3D motion-music correlation. This metric evaluates synchronization between motion and music in 2D scenarios, demonstrating the effectiveness of our approach in producing rhythmically aligned dance videos.
@@
Our contribution can be summarized as follows:
  • We introduce a novel approach capable of directly generating dance videos from music, significantly enhancing the intuitiveness of the generated content and its applicability in real-world settings.
  • Our research broadens the horizon of conditional image-to-video generation by exploring music as a novel conditional input. We investigate the extraction of meaningful information from music through the integration of large-scale foundation models and established signal-processing methodologies.
  • We establish a baseline and propose a beat alignment module for this innovative task, supported by extensive evaluations and the introduction of a new evaluation metric to assess motion-music alignment.
@@ 2 Related Works
Generating dance sequences from music uniquely intersects motion synthesis [1, 4, 12, 14] and music interpretation [58, 30, 59], aiming to create choreographed movements synchronized with input music. This extends beyond traditional motion synthesis, as choreographed movements are complex to animate. Early research [43, 69, 9] focused on producing 2D dance sequences due to the availability of online dance videos and advances in 2D human pose estimation [5]. However, 2D predictions lack expressiveness and applicability, prompting a shift to 3D dance generation. Recent methods leverage LSTMs [49, 2, 20, 61], GANs [25, 48, 10, 21], and Transformers [17, 27, 47, 35, 64] for 3D motion generation. The AIST++ dataset [27], a rich compilation of 3D motion data, has significantly advanced the field. Notable models include FACT [27], a full-attention-based cross-modal transformer using sequence-to-sequence learning for lifelike 3D dance sequences, and Bailando [47], which combines a pose VQ-VAE with a Motion GPT for temporal coherence via actor-critic learning. While existing research focuses on generating motion sequences from music, Ren et al. [37] synthesize videos using these sequences. Our work bypasses intermediary joint sequence generation, directly creating dance videos from music.
@@ 3 Method
Our methodology aims to generate latent optical flows guided by musical inputs. Recent works [31, 44, 52, 53, 32] in motion transfer have shown the effectiveness of using latent optical flow for warping images. This approach is more resource-efficient, requiring less computational power and time compared to high-dimensional pixel or latent feature spaces. Diffusion models are chosen for their exceptional quality and robust controllability. Our methodology begins by training an auto-encoder to discern optical flow between video frames. This trained auto-encoder then aids in training the diffusion model to generate latent flows. We extract musical information using CLAP to encode the music. Our enhanced model also incorporates beat information for better representation.
Figure 3: Training of latent flow auto-encoder. The flow predictor learns to estimate the latent flow $f$ and occlusion map $m$ between the reference frame $x_{ref}$ and driving frame $x_{dri}$. The image encoder encodes $x_{ref}$ into a latent representation $z$, $f$ and $m$ are utilized to manipulate $z$ into $\tilde{z}$ which is then decoded by an image decoder to generate a output image $\hat{x}_{out}$. The objective of the training is to minimize the disparity between $x_{dri}$ and $\hat{x}_{out}$.
@@ 3.1 Latent Flow Estimation
The objective of this phase is the training of an auto-encoder capable of accurately capturing and modeling motion between video frames within a latent space. As illustrated in Figure 3, the architectural components of the model comprise an image encoder, a flow predictor, and an image decoder. During the training process, two frames are randomly selected from a single video to serve as a reference frame $x_{ref}$ and a driving frame $x_{dri}$, both sharing dimensions of $H_{x}\times W_{x}\times 3$. The image encoder processes $x_{ref}$, converting it into a compact latent representation, $z$, with dimensions $H_{z}\times W_{z}\times C_{z}$. Subsequently, similar to [44, 53], the flow predictor, receiving both $x_{ref}$ and $x_{dri}$ as inputs, computes the backward latent optical flow, $f$, and an occlusion map, $m$, to denote the transformation between these frames. The estimated latent flow, $f$, maintains the spatial dimensions of $z$ and incorporates two channels to articulate the horizontal and vertical displacement across frames. Using backward flow estimation, the warping of $z$ using $f$ is realized through an efficient differentiable bilinear sampling [18]. Concurrently, the occlusion map, sized $H_{z}\times W_{z}\times 1$, facilitates the reconstruction of areas within $z$ that become obscured or revealed due to movement. Values within this map range from 0 to 1, where 1 signifies unoccluded regions, and 0 denotes complete occlusion. The warped latent map, $\tilde{z}$, is derived through the following equation:
@@
Figure 4: Inference process of DabFusion. DabFusion incorporates noise input along with image embedding $z_{0}$ and music embedding $e$ as conditions. Following the denoising stage of the diffusion model, we obtain $h_{1}^{N}$, comprising a concatenated sequence of latent flow and corresponding occlusion maps. $h_{1}^{N}$ is utilized to transform $z_{0}$ into a new sequence of latent maps, denoted as $\tilde{z}_{1}^{N}$, which is subsequently decoded to produce an image sequence.
@@ 3.2 Latent Flow Generation
Our approach utilizes a 3D U-Net [7] as the denoising mode. And the training of the model is in conjunction with the trained image encoder and flow predictor. Given an input video $x_{0}^{N}=\{x_{0},x_{1},...,x_{N}\}$ and its corresponding music $m$. We first use the flow predictor to estimate latent flow sequence $f_{1}^{N}=\{f_{1},...,f_{N}\}$ and occlusion map sequence $m_{1}^{N}=\{m_{1},...,m_{N}\}$ between starting frame $x_{0}$ and remaining frames $\{x_{1},...,x_{N}\}$. The size of $f_{1}^{N}$ is $N\times H_{z}\times W_{z}\times 2$ and the size of $m_{1}^{N}$ is $N\times H_{z}\times W_{z}\times 1$. Subsequently, we concatenate these two sequences along the last dimension to get $a_{0}=[f_{1}^{N},m_{1}^{N}]$ which has the size $N\times H_{z}\times W_{z}\times 3$. Then, we perform diffusion process which gradually add 3D Gaussian noise to $a_{0}$ to map it to a a standard Gaussian noise. This is then reversed, using the image encoder to map the initial frame $x_{0}$ to latent space $z_{0}$ and incorporating music embeddings e for conditioning the denoising model. We will discuss how to encode music in Sec. 3.3. The updated loss function, accounting for the conditioning on $z_{0}$ and e, is given by:
@@
The process of inference is shown in Figure 4 , given an initial image $x_{0}$ and a piece of music $m$, the image is first encoded into a latent space representation $z_{0}$, while the music is transformed into an embedding $e$. Subsequently, a volume of randomly sampled Gaussian noise, with dimensions $N\times H_{z}\times W_{z}\times 3$, undergoes a gradual denoising process by the U-Net. This process yields $h_{1}^{N}$ which is the concatenation of a latent flow sequence $f_{1}^{N}$ and corresponding occlusion map sequence $m_{1}^{N}$. $z_{0}$ is then warped by each latent flow in the latent flow sequence and corresponding occlusion map in the occlusion map sequence to generate a new latent map sequence $\tilde{z}_{1}^{N}$. These maps are sequentially fed into the image decoder, which synthesizes the frames of the new video.
3.3 Music Encoding
Extracting meaningful information from music is crucial to our task. We use CLAP, developed by Wu et al. [59], for music encoding. CLAP is a state-of-the-art audio representation model trained on a large dataset of text-audio pairs, learning a joint text-audio embedding space through contrastive learning. Our choice of CLAP is strategic, aiming to decode dance style information from musical cues using text as a bridge for music-to-dance style translation. While CLAP excels at encoding musical style, it struggles with detecting rhythmic patterns. To address this, we integrate beat information extracted via Librosa with CLAP’s output, creating a composite music embedding. This enhanced approach improves synchronization between motion and beat compared to our baseline model.
4 Experiments
@@ 4.1 Dataset and Implementation Details
We organize our dataset by music pieces, allocating 10,564 videos and 50 music tracks to the training set, and 2,106 videos with 10 tracks to the test set. Each video is segmented into frames, resized to $128\times 128$ pixels. In the latent flow estimation phase, we adopt the architecture from [19] for both our image encoder and decoder. The flow predictor follows [44]. The model is trained over 150 epochs with a batch size of 100, using the Adam optimizer [23], starting with a learning rate of $2\times 10^{-4}$, decreased by a factor of 0.1 after epochs 60, 90, and 120. For the denoising model’s training, we employ a conditional 3D U-Net architecture from [13], featuring four down-sampling and up-sampling 3D convolutional blocks. Time step $t$ is encoded by the sinusoidal position embedding [51]. The training spanned 250 epochs, starting with the same initial learning rate as the first stage, and adjusted at epochs 100, 150, and 200. We set $T$ = 1000 and apply a cosine noise schedule [33] and dynamic thresholding [41] at the 90% during sampling.
Figure 5: Video quality comparison between our models. We use the same starting image and music piece to generate videos with both our models, and select three frames from the same position.

### v2

@@ Abstract
Automated choreography advances by generating dance from music. Current methods create skeleton keypoint sequences, not full dance videos, and cannot make specific individuals dance, limiting their real-world use. These methods also need precise keypoint annotations, making data collection difficult and restricting the use of self-made video datasets. To overcome these challenges, we introduce a novel task: generating dance videos directly from images of individuals guided by music. This task enables the dance generation of specific individuals without requiring keypoint annotations, making it more versatile and applicable to various situations. Our solution, the Dance Any Beat Diffusion model (DabFusion), utilizes a reference image and a music piece to generate dance videos featuring various dance types and choreographies. The music is analyzed by our specially designed music encoder, which identifies essential features including dance style, movement, and rhythm. DabFusion excels in generating dance videos not only for individuals in the training dataset but also for any previously unseen person. This versatility stems from its approach of generating latent optical flow, which contains all necessary motion information to animate any person in the image. We evaluate DabFusion’s performance using the AIST++ dataset, focusing on video quality, audio-video synchronization, and motion-music alignment. We propose a 2D Motion-Music Alignment Score (2D-MM Align), which builds on the Beat Alignment Score to more effectively evaluate motion-music alignment for this new task. Experiments show that our DabFusion establishes a solid baseline for this innovative task. Video results can be found on our project page: https://DabFusion.github.io.
@@ 1 Introduction
Generating dance from music has great potential for automating choreography and enhancing various applications. Current methods [25, 26, 46, 34] primarily produce keypoint sequences which, while effective, often yield outputs that are not as user-friendly and cannot animate specific individuals, limiting their use in personalized scenarios. These methods also rely heavily on precisely annotated keypoints, complicating data collection and limiting their application to non-standard or self-made video datasets. To address these limitations, we introduce a new and innovative task: generating dance videos directly from images of individuals, guided by accompanying music. This allows for personalized dance videos featuring specific people without needing keypoint annotations, making it more versatile for different scenarios. Our solution, named the Dance Any Beat Diffusion model (DabFusion), takes a reference image paired with a selected piece of music as input to create dance videos featuring various dance types and choreographies. For extracting information from music, we develop a novel music encoder that captures essential features including dance style, movement, and rhythm, which are critical for this task. To accurately extract dance styles, we fine-tune the CLAP [58] model, a text-audio pretraining foundation model, on the AIST++ [26] dataset, which serves as the training dataset for our task. For movement information, we fine-tune the Wav2CLIP [57] model, a visual-audio pretraining foundation model, on the AIST++ dataset. The synchronization of dance movements with the music’s rhythm is essential for realistic dance generation. To achieve this, we employ Librosa [29], a tool for audio signal analysis, to extract beat information from the music. By integrating these elements, our music encoder can generate a comprehensive representation of the music.
DabFusion effectively generates dance videos not just for individuals included in the AIST++ dataset, but also for people it has never encountered before. As shown in Fig. 1, we animate previously unseen individuals by merging their images with a background from the AIST++ dataset, bringing them to life through dance. DabFusion can choreograph dance sequences for previously unseen individuals to various music pieces, producing distinctly different videos. This diversity in output highlights the robust information extraction capability of our music encoder. When generating dance videos using the AIST++ dataset, as illustrated in Fig. 2, DabFusion excels in producing varied dance styles with different dancers from multiple perspectives and diverse initial poses. The capability of DabFusion to generate various dance forms enables it to serve a broad audience, ranging from professional choreographers seeking inspiration to casual users wanting to see themselves or others dance to their favorite tunes. Additionally, this technology can be utilized in entertainment applications, such as music videos and virtual concerts, where personalized and accurately synchronized dance performances can greatly enhance the viewer experience. The versatility of DabFusion arises from its method of generating latent optical flow, which captures all essential motion information needed to animate any person in the image. This technique is widely adopted in the field of image-to-video (I2V) generation [53, 38, 10, 5, 64, 54, 37, 44, 43, 67]. Inspired by recent I2V works [31, 43, 56], our model’s training process is structured into two phases. First, a latent flow auto-encoder is trained unsupervisedly to estimate the latent optical flow between reference and driving frames in a video, aiding in warping the reference frame to generate movement. Then, a U-Net based diffusion model generates latent flows guided by music and a starting image. We conduct a thorough evaluation of DabFusion’s performance using the AIST++ dataset, with particular focus on video quality, audio-video synchronization, and the alignment between motion and music. To more effectively assess motion-music alignment—a crucial aspect of this new task—we introduce a 2D motion-music alignment (2D-MM Align) score, inspired by the Beat Alignment Score [26] for 3D motion-music correlation. This metric evaluates synchronization between motion and music in 2D scenarios, demonstrating the effectiveness of our approach in producing rhythmically aligned dance videos. Our experimental results demonstrate that DabFusion establishes a robust baseline for this innovative task, indicating its potential as a pioneering tool in the field of dance video synthesis.
Our contributions can be summarized as follows:
  • We introduce a novel task that directly generates dance videos from music, significantly enhancing the intuitiveness of the generated content and its applicability in real-world settings.
  • We propose DabFusion, a solution for the new task. DabFusion is capable of creating dance sequences not only for individuals present in the training dataset but also for person it has never previously encountered.
  • We propose a novel music encoder designed to generate a comprehensive representation of music, capturing key elements including dance style, movement, and rhythm.
  • We establish a solid baseline for the new task, supported by extensive evaluations on the AIST++ dataset.
@@ 2 Related Works
Generating dance sequences from music uniquely intersects motion synthesis [1, 3, 11, 13] and music interpretation [57, 29, 58], aiming to create choreographed movements synchronized with input music. This extends beyond traditional motion synthesis, as choreographed movements are complex to animate. Early research [42, 68, 8] focused on producing 2D dance sequences due to the availability of online dance videos and advances in 2D human pose estimation [4]. However, 2D predictions lack expressiveness and applicability, prompting a shift to 3D dance generation. Recent methods leverage LSTMs [48, 19, 60], GANs [24, 47, 9, 20], and Transformers [16, 26, 46, 34, 63] for 3D motion generation. The AIST++ dataset [26], a rich compilation of 3D motion data, has significantly advanced the field. Notable models include FACT [26], a full-attention-based cross-modal transformer using sequence-to-sequence learning for lifelike 3D dance sequences, and Bailando [46], which combines a pose VQ-VAE with a Motion GPT for temporal coherence via actor-critic learning. While existing research focuses on generating motion sequences from music, Ren et al. [36] synthesizes videos using these sequences. Our work bypasses intermediary joint sequence generation, directly creating dance videos from music.
@@
Figure 3: Overview of DabFusion. DabFusion incorporates noise input along with image embedding $z_{0}$ and music embedding $e$ as conditions. Following the denoising stage of the diffusion model, we obtain $h_{1}^{N}$, comprising a concatenated sequence of latent flow and corresponding occlusion maps. $h_{1}^{N}$ is utilized to transform $z_{0}$ into a new sequence of latent maps, denoted as $\tilde{z}_{1}^{N}$, which is subsequently decoded to produce an image sequence.
@@ 3 Method
Our methodology generates latent optical flows driven by musical inputs, with each music piece encoded by our specially designed music encoder. Music encoding is detailed in Sec. 3.1. Recent works [30, 43, 51, 52, 31] in motion transfer have shown the effectiveness of using latent optical flow for warping images. This approach is more resource-efficient, requiring less computational power and time compared to high-dimensional pixel or latent feature spaces. Our methodology begins by training an auto-encoder to discern optical flow between video frames, we discuss this in details in Sec. 3.2. This trained auto-encoder then aids in training the diffusion model to generate latent flows. The training of the diffusion model is detailed in Sec. 3.3. The overview of DabFusion is shown in Fig. 3 , given an initial image $x_{0}$ and a piece of music $m$, the image is first encoded into a latent space representation $z_{0}$, while the music is transformed into an embedding $e$. Subsequently, a volume of randomly sampled Gaussian noise, with dimensions $N\times H_{z}\times W_{z}\times 3$, undergoes a gradual denoising process by the U-Net. This process yields $h_{1}^{N}$ which is the concatenation of a latent flow sequence $f_{1}^{N}$ and corresponding occlusion map sequence $m_{1}^{N}$. $z_{0}$ is then warped by each latent flow in the latent flow sequence and corresponding occlusion map in the occlusion map sequence to generate a new latent map sequence $\tilde{z}_{1}^{N}$. These maps are sequentially fed into the image decoder, which synthesizes the frames of the new video.
3.1 Music Encoding
Extracting meaningful information from music is essential for generating dance videos, where dance type, dancer movement, and rhythm play critical roles. For dance type information, we utilize the CLAP model, a pipeline for contrastive language-audio pretraining that develops audio representations by integrating audio data with natural language descriptions. We fine-tune CLAP to enhance its capture of dance style information. Specifically, we freeze the weights of the audio and text encoders in CLAP and introduce multi-layer perceptrons as adapter layers. These layers are trained using data from the AIST++ dataset. For movement information, we employ Wav2CLIP, initially trained on audio-visual datasets. We apply a similar fine-tuning approach as with CLAP, using the AIST++ dataset to adapt Wav2CLIP for our specific requirements. Regarding rhythm, which is crucial for aligning dance movements with the music beat, we use Librosa, a tool for audio signal analysis, to extract beat information. We then concatenate these diverse sources of information to create a comprehensive musical representation tailored for dance video generation.
Figure 4: Training of latent flow auto-encoder. The flow predictor learns to estimate the latent flow $f$ and occlusion map $m$ between the reference frame $x_{ref}$ and the driving frame $x_{dri}$. The image encoder encodes $x_{ref}$ into a latent representation $z$, $f$ and $m$ are utilized to manipulate $z$ into $\tilde{z}$ which is then decoded by an image decoder to generate an output image $\hat{x}_{out}$. The objective of the training is to minimize the disparity between $x_{dri}$ and $\hat{x}_{out}$.
@@ 3.2 Latent Flow Estimation
The objective of this phase is the training of an auto-encoder capable of accurately capturing and modeling motion between video frames within a latent space. As illustrated in Fig. 4, the architectural components of the model comprise an image encoder, a flow predictor, and an image decoder. During the training process, two frames are randomly selected from a single video to serve as a reference frame $x_{ref}$ and a driving frame $x_{dri}$, both sharing dimensions of $H_{x}\times W_{x}\times 3$. The image encoder processes $x_{ref}$, converting it into a compact latent representation, $z$, with dimensions $H_{z}\times W_{z}\times C_{z}$. Subsequently, similar to [43, 52], the flow predictor, receiving both $x_{ref}$ and $x_{dri}$ as inputs, computes the backward latent optical flow, $f$, and an occlusion map, $m$, to denote the transformation between these frames. The estimated latent flow, $f$, maintains the spatial dimensions of $z$ and incorporates two channels to articulate the horizontal and vertical displacement across frames. Using backward flow estimation, the warping of $z$ using $f$ is realized through an efficient differentiable bilinear sampling [17]. Concurrently, the occlusion map, sized $H_{z}\times W_{z}\times 1$, facilitates the reconstruction of areas within $z$ that become obscured or revealed due to movement. Values within this map range from 0 to 1, where 1 signifies unoccluded regions, and 0 denotes complete occlusion. The warped latent map, $\tilde{z}$, is derived through the following equation:
@@ 3.3 Latent Flow Generation
Our approach utilizes a 3D U-Net [6] as the denoising mode. And the training of the model is in conjunction with the trained image encoder and flow predictor. Given an input video $x_{0}^{N}=\{x_{0},x_{1},...,x_{N}\}$ and its corresponding music $m$, we first use the flow predictor to estimate latent flow sequence $f_{1}^{N}=\{f_{1},...,f_{N}\}$ and occlusion map sequence $m_{1}^{N}=\{m_{1},...,m_{N}\}$ between starting frame $x_{0}$ and remaining frames $\{x_{1},...,x_{N}\}$. The size of $f_{1}^{N}$ is $N\times H_{z}\times W_{z}\times 2$ and the size of $m_{1}^{N}$ is $N\times H_{z}\times W_{z}\times 1$. Subsequently, we concatenate these two sequences along the last dimension to get $a_{0}=[f_{1}^{N},m_{1}^{N}]$ which has the size $N\times H_{z}\times W_{z}\times 3$. Then, we perform diffusion process which gradually adds 3D Gaussian noise to $a_{0}$ to map it to a a standard Gaussian noise. This is then reversed, using the image encoder to map the initial frame $x_{0}$ to latent space $z_{0}$ and incorporating music embeddings $e$ for conditioning the denoising model. The updated loss function, accounting for the conditioning on $z_{0}$ and $e$, is given by:
@@
Figure 5: Comparison of video quality between ground-truth videos and those generated by DabFusion. We use the same starting image and music piece to generate videos with our model, and select three frames from the same position.
4 Experiments and Results
@@ 4.1 Dataset and Implementation Details
We organize our dataset by music pieces, allocating 10,564 videos and 50 music tracks to the training set, and 2,106 videos with 10 tracks to the test set. Each video is segmented into frames, resized to $128\times 128$ pixels. In the latent flow estimation phase, we adopt the architecture from [18] for both our image encoder and decoder. The flow predictor follows [43]. The model is trained over 150 epochs with a batch size of 100, using the Adam optimizer [22], starting with a learning rate of $2\times 10^{-4}$, decreased by a factor of 0.1 after epochs 60, 90, and 120. For the denoising model’s training, we employ a conditional 3D U-Net architecture from [12], featuring four down-sampling and up-sampling 3D convolutional blocks. The time step $t$ is encoded by the sinusoidal position embedding [50]. The training spanned 250 epochs, starting with the same initial learning rate as the first stage, and adjusted at epochs 100, 150, and 200. We set $T$ = 1000 and apply a cosine noise schedule [32] and dynamic thresholding [40] at the 90% during sampling.